# Biological computation through recurrence

### 07 Feb 2024 v1

@@ IV Evolutionary aspects of recurrence-based computations
When RC was initially introduced, it was shown mathematically that the reservoir needs to have two fundamental attributes for the proper implementation of the paradigm [14]. Firstly, reservoir dynamics must be able to separate input features that are significantly different for a range of tasks. Secondly, these same dynamics should be able to generalize to new examples, thus effectively mapping similar inputs to proximate regions within the dynamical reservoir space.
Legenstein and Maass [52] analyzed how reservoir dynamics contributes to those two requirements, showing that separability and generalization are anticorrelated with each other: while recurrent networks with simple attractor dynamics have substantial generalization capabilities (because large classes of noisy inputs result in converging dynamics), they exhibit low separability. On the other hand, recurrent networks with chaotic dynamics have large separability but very low generalization capabilities (because chaotic systems are highly sensitive to initial conditions).
Consequently, to fulfill simultaneously the conditions of separability and generalization, recurrent networks need to operate in between ordered and chaotic dynamics, i.e. near the so-called edge of chaos. This refers to a critical state between ordered dynamics (where disturbances quickly die out into attractors) and chaotic dynamics (where disturbances are amplified). This property is particularly interesting given the evidence in the literature that cortical circuits are tuned to criticality [53, 54, 55].
In this sense, several authors have used hallmark indicators of criticality to design optimal reservoirs with enhanced performance. The measures employed include Lyapunov exponents [52, 56, 57, 58, 59] and Fisher information [60] of reservoir dynamics. Taken together, these studies highlight the advantages of criticality for information processing, and suggest powerful evolutionary constraints that should bring naturally occurring RC-based systems towards the edge of chaos.
@@ V Final remarks
In spite of their conceptual appeal, validating experimentally the role of recurrence-based computations in biological systems is still challenging. For example, while the existence of recurrent architectures in cortical brain regions has been examined in many works, as described above, studies linking systematically the dynamics of recurrent cortical networks with time-dependent tasks are still needed. To address this issue, cognitive systems should be evaluated simultaneously from the structural, dynamical, and functional viewpoints.
@@ Acknowledgments
This work was supported by project PID2021-127311NB-I00, financed by the Spanish Ministry of Science and Innovation, the Spanish State Research Agency and FEDER (MICIN/AEI/10.13039/ 501100011033/FEDER). Financial support was also provided by the Maria de Maeztu Programme for Units of Excellence in R&D (Spanish State Research Agency, project CEX2018-000792-M), and by the ICREA Academia programme. M.S.V. is supported by a PhD fellowship from the Agència de Gestió d’Ajuts Universitaris i de Recerca (AGAUR) from the Generalitat de Catalunya (grant 2021-FI-B-00408).

### 25 Feb 2024 v2

@@ IV Evolutionary aspects of recurrence-based computations
When RC was initially introduced, it was shown mathematically that the reservoir needs to have two fundamental attributes for the proper implementation of the paradigm [14]. Firstly, the dynamics of the reservoir needs to be sufficiently high-dimensional to effectively separate training inputs across a variety of tasks. Secondly, the system should be able to generalize its classification abilities to new examples. This would entail mapping similar inputs to nearby regions within the dynamical reservoir’s phase space.
Legenstein and Maass [51] analyzed how reservoir dynamics contributes to those two requirements, showing that separability and generalization are anticorrelated with each other. Recurrent networks with simple attractor dynamics, for instance, have substantial generalization capabilities (because even under noise, many inputs give rise to similar dynamics in the reservoir’s phase space), they exhibit low separability. When recurrent networks operate in the chaotic regime, in contrast, the dynamics becomes highly sensitive to initial conditions, which strongly impairs generalization, while separability remains large due to the nonlinear and high-dimensional character of the system.
Consequently, to fulfill simultaneously the conditions of separability and generalization, recurrent networks need to operate near the so-called edge of chaos, at the boundary between order and chaos. This critical region represents a compromise between attractor dynamics, in which perturbations (even relatively large ones) are washed away as the network approaches equilibrium, providing generalization, and chaotic dynamics, where perturbations (even very small ones) are amplified, providing separability. This requirement is consistent with the large amount of evidence reported during the past two decades showing that the brain operates near criticality[52, 53, 54].
Following this line of reasoning, criticality has been used in recent years as a global constraint in the design of optimal reservoirs, employing quantifiers such as Lyapunov exponents [51, 55, 56, 57, 58] and Fisher information [59] of reservoir dynamics. Taken together, these studies highlight the advantages of criticality for information processing, and suggest powerful constraints that could have played a relevant role in evolution of the computational capabilities of recurrent biological networks.
@@ V Final remarks
In spite of their conceptual appeal, validating experimentally the role of recurrence-based computations in biological systems is still challenging. For example, while the existence of recurrent architectures in cortical brain regions has been thoroughly studied, as described above, research linking systematically the dynamics of recurrent cortical networks with time-dependent tasks is still needed. This will require a comprehensive analysis of cognitive systems at the structural, dynamical, and functional levels.
@@ Acknowledgments
This work was supported by project PID2021-127311NB-I00 / MICIN / AEI / 10.13039 / 501100 011033 / FEDER-UE, financed by the Spanish Ministry of Science and Innovation, the Spanish State Research Agency and the European Regional Development Fund (FEDER). Financial support was also provided by the Maria de Maeztu Programme for Units of Excellence in R&D (Spanish State Research Agency, project CEX2018-000792-M), and by the ICREA Academia programme. M.S.V. is supported by a PhD fellowship from the Agència de Gestió d’Ajuts Universitaris i de Recerca (AGAUR) from the Generalitat de Catalunya (grant 2021-FI-B-00408).